OpenStudy (anonymous):

show that sin 18 is the soluction to f(x)=8x^3-4x+1

OpenStudy (anonymous):

without the use of a calfram

OpenStudy (anonymous):

less of a calculus more of a trig ?

OpenStudy (anonymous):

\[Hint :4\sin18*\cos36=1\]

OpenStudy (anonymous):

\[x=18\]\[5x=90\]\[3x=90-2x\]\[\sin 3x=\cos 2x\]\[\sin x(3-4\sin^2 x)=\cos^2 x-\sin^2 x=1-2\sin^2 x\]\[y=\sin x\]\[y(3-4y^2)=1-2y^2\] where did i do wrong?

OpenStudy (anonymous):

\[4y^3-2y^2-3y+1=0\]

OpenStudy (anonymous):

urs is correct too :)

OpenStudy (anonymous):

anyway...i learned this method from @Callisto

OpenStudy (anonymous):

jonask how u get that ???\[4 \sin 18 \cos 36=1\]

OpenStudy (anonymous):

you used 3 instead of 8

OpenStudy (anonymous):

\[4\cos36=\frac{ 1 }{ \sin18 }\]

OpenStudy (anonymous):

it was an identity given and these is one of those hence question s do you think i should post the identity

OpenStudy (anonymous):

easy solution will be veryfying by substitution \[8\sin^318-4\sin18+1\] \[4\sin18(2\sin^218-1)+1\] \[4\sin18(-\cos36)+1\] \[\frac{ 1 }{ \cos36 }(-\cos36)+1=-1+1=0\]

OpenStudy (anonymous):

what method is sir @Callisto

OpenStudy (anonymous):

what method did you show @mukushla

OpenStudy (callisto):

Referring to your question, actually, part b gives you a big hint. \[1 = 4cos36sin18\]\[1= 4(1-2sin^218)sin18\]\[1= 4sin18-8sin^318\]Rearrange the terms and you'll get the equation in the form \[8x^2 - 4x + 1 =0\] I'm still thinking why the method introduced by @mukushla doesn't work.

OpenStudy (anonymous):

wow thats relly an answer in disguise more than a Hint...well i see @mukushla printed sinx(3-4sinx) wasn't it suppsedto be sinx(8-4sinx)

OpenStudy (callisto):

Hmm.. Usually for questions with part a, b, c and so on, the previous parts help you do the rest of the question. As for mukushla's workings, he expanded sin3x to get that. sin3x = sin(2x+x) = sin2x cosx + cos2x sinx = (2sinx cosx) cosx + (1-2sin^2 x) sinx = 2sinx cos^2x + (1-2sin^2 x) sinx = 2sinx - 2sin^3 x + sinx - 2sin^3 x = 3sinx - 4sin^3 x = sinx (3-4sin^2 x)

OpenStudy (anonymous):

i guess the method is fine as well but you get a different answer

OpenStudy (anonymous):

answer that looks different but is the same i mean

OpenStudy (callisto):

That method would be fine, the only problem is how to get the equation in the form of 8x^3 - 4x +1 =0 :S

OpenStudy (anonymous):

we need to show that 4y^3 + 2y^2 - y = 0 which is a problem by itself

OpenStudy (callisto):

Of course that's the problem, which we need to solve.

OpenStudy (anonymous):

i meant that 4y^3 + 2y^2 - y this is what is different between the answers

OpenStudy (callisto):

I understand. But we have to show that 4y^3+2y^2-y =0.

OpenStudy (anonymous):

the other problem is that sin 18 can be a solution of many other curves and it happens that @mukushla is one of them

OpenStudy (callisto):

@Jonask Actually, you just have to show that sin 18 is the solution to f(x)=8x^3-4x+1. It doesn't matter if it is a solution to other equation. And using the result of part b, you can do it. (well, I just show you). We're thinking if there are no hints provided, which means we don't know the result of part b, and using the method introduced by mukushla here. How can we get the answer. Now, we get to the point which we have to prove 4y^3+2y^2-y = 0.

OpenStudy (anonymous):

im giving up .. this method is wonderful anyway ive learned something

OpenStudy (anonymous):

so we want to prove that when x=sin18

OpenStudy (callisto):

Yes. We set x=18 and y = sin18 in mukushla's work.

OpenStudy (anonymous):

i notice that we changed \[4y^3-2y^2-3y+1=0\rightarrow 4y^3+2y^2-y=0\]

OpenStudy (anonymous):

which also works for sin18

OpenStudy (callisto):

Not really. Now, we have \(4y^3-2y^2-3y+1=0\) and we want to show that \(8y^3 - 4y +1 =0\) So, add 4y^3 and subtract 2y^2 and y from the left of \(4y^3-2y^2-3y+1=0\), and at the same time, add 4y^3, 2y^2 and subtract y from the right of \(4y^3-2y^2-3y+1=0\). Then we get. \(4y^3-2y^2-3y+1+4y^3+2y^2 -y= 4y^3 + 2y^2 -y\) \(8y^3-4y+1= 4y^3 + 2y^2 -y\) Compare \(8y^3-4y+1= 4y^3 + 2y^2 -y\) (something we have now) with \(8y^3 - 4y +1 =0\) (something we need) We need to prove that 4y^3 + 2y^2 -y =0 (by looking at the right side of both equations), where y = sin18

OpenStudy (callisto):

@Jonask Btw, actually, you can solve you problem now. You don't have to solve it using this method.. :S

OpenStudy (anonymous):

thanks,this is interesting thanks again

OpenStudy (callisto):

Welcome!~ This is very interesting indeed! :)